OpenStudy (anonymous):

An equation is shown below: 4(2x − 3) = 1 Which of the following correctly shows the first two steps to solve this equation?

OpenStudy (anonymous):

A. Step 1: 6x − 7 = 1; Step 2: 6x = 8 B. Step 1: 6x + 1 = 1; Step 2: 6x = 0 C. Step 1: 8x − 12 = 1; Step 2: 8x = 13 D. Step 1: 8x − 3 = 1; Step 2: 8x = 4

OpenStudy (f_jayyy):

C

OpenStudy (f_jayyy):

because you would first need to multiply 4 by the 2x and 3 ....so you get 8x-12

OpenStudy (f_jayyy):

8x-12=1 8x=1+12 8x=13 so the answer is C

OpenStudy (anonymous):

Thank you it was correct :)

OpenStudy (f_jayyy):

np